OpenStudy (anonymous):

the kentucky lottery is worth $100,000. to win you have to select the 5 winning numbers from 1-50. how many different combinations are possible?

OpenStudy (lgbasallote):

i believe the formula would be \[\large \frac{50!}{5! 45!}\] feel free to correct me though @apoorvk

OpenStudy (anonymous):

where did you get the 45?

OpenStudy (lgbasallote):

50 - 5 i believe

Parth (parthkohli):

Yes, it'd be 50c5 \(\Large \color{purple}{ 50! \over (50 - 5)!*5! }\)

Parth (parthkohli):

Get it?

OpenStudy (amorfide):



Parth (parthkohli):

nCr is represented as: \(\Large \color{purple}{ {n! \over (n - r)!*r!} }\)

OpenStudy (amorfide):

n= the amount of numbers =50 r= the numbers you want so 5 50!/(50-5)!5!

OpenStudy (apoorvk):

no @lgbasallote I won't need to correct you. For choosing 'r' things out 'n', (arrangement not being considered), no. of possible combinations 'N': \[\large N = ^nC_r = \frac{n!}{r!(n-r)!}\]

OpenStudy (lgbasallote):

heh i cant believe i got the formula right..i only context clues-ed from amistre's work earlier \m/

Parth (parthkohli):

If you have to choose 5 numbers from 50, then it'll be 50 choose 5 aka 50c5

OpenStudy (lgbasallote):

can you go from there now @catgirl17

Parth (parthkohli):

Considering that you know what factorials are lolz

OpenStudy (amorfide):

oh god please know what factorials are...

OpenStudy (lgbasallote):

it's gonna be a long post slash night if fno

Parth (parthkohli):

\(\Large \color{purple}{ n! = n * (n - 1) * (n - 2) * (n - 3) ........1 }\)

OpenStudy (anonymous):

after 50!/45!x5! do i just expand it out

Parth (parthkohli):

Just having some fun

OpenStudy (lgbasallote):

yes you do @catgirl17

Parth (parthkohli):

And then you cancel to make it simpler.

OpenStudy (lgbasallote):

and if i were you get rid of x...you might mistake it for a variable

OpenStudy (anonymous):

thank you @lgbasallote

Parth (parthkohli):

Use * instead of x lol

OpenStudy (lgbasallote):

\[\huge \color{maroon}{\mathtt{\text{<tips hat>}}}\]

Parth (parthkohli):

* doesn't only represent the thing that twinkles. lol

OpenStudy (lgbasallote):

or use ( )( ) yeah...they're more subtle

Parth (parthkohli):

\(\Large \color{purple}{\mathbb{Still} \text{ doesn't get the medal. xP} }\)

OpenStudy (anonymous):

when expanding the 50! out do i have to go all the way to 5! ?

OpenStudy (amorfide):

50! =50x49x48x47......x1

Parth (parthkohli):

wait wait... n! OVER r! = (n - r)!

OpenStudy (amorfide):

50!/45! cancels out to 50x49x48x47x46 divide that by 5!

OpenStudy (lgbasallote):

just until 45! it will cancel out anyway... \[\large \frac{50! 49! 48! 47! 46! 45!}{45! 5!}\] see? 45! will just cancel out

OpenStudy (amorfide):

why have you got all those factorials on top?!

Parth (parthkohli):

Are you kdding me? 50! is not 50! * 49!........

OpenStudy (lgbasallote):

fine get rid of the ! :P

OpenStudy (amorfide):